# Token-Efficient RL for LLM Reasoning

### v2

@@ Abstract
We propose reinforcement learning (RL) strategies tailored for reasoning in large language models (LLMs) under strict memory and compute limits, with a particular focus on compatibility with LoRA fine-tuning. Rather than relying on full-sequence updates or separate critic networks, we design critic-free methods that operate on a small, informative subset of output tokens to reduce memory usage and stabilize training. We introduce S-GRPO, a stochastic variant of Group Relative Policy Optimization, and T-SPMO, a token-level prefix matching approach for fine-grained credit assignment. Applied to Qwen2-1.5B, our methods raise accuracy on the SVAMP benchmark from 46% to over 70% and show strong performance on multi-digit multiplication. Surprisingly, full-token GRPO under LoRA fails to improve over the base model, suggesting that selective token-level optimization may act as an implicit regularizer in low-parameter training regimes.
@@ 1 Introduction
In this work, we design and evaluate RL algorithms that are explicitly tailored to low compute constraints: Partial parameter updates, limited GPU capacity, and the need to mitigate overfitting in low-capacity adaptation settings. First, we introduce Stochastic-GRPO (S-GRPO), a lightweight variant of GRPO that samples tokens from output trajectories to contribute to the loss. Second, we propose Token-Specific Prefix Matching Optimization (T-SPMO), which assigns credit at a token granularity.
To evaluate these methods, we fine-tune the open-weight Qwen2-1.5B model (Yang et al., 2024) using LoRA (Hu et al., 2021) on two reasoning benchmarks: the SVAMP dataset (Patel et al., 2021) and a multi-digit multiplication task. Both S-GRPO and T-SPMO significantly raise test-set accuracy on both benchmarks. We also establish a full-token GRPO baseline under LoRA, which surprisingly did not improve performance from the base model. These results suggest that carefully designed, memory-efficient RL algorithms can significantly improve LLM reasoning under compute constraints—even outperforming exhaustive full-token optimization strategies.
@@ 2 Related Work
2.1 RL for Large Language Models
RL has been widely used to align language model outputs with user preferences or task-specific objectives. Proximal Policy Optimization (PPO) (Schulman et al., 2017) is a commonly used algorithm in RL from human feedback (RLHF), most notably applied in training models like InstructGPT (Ouyang et al., 2022). However, PPO requires training a separate value network (critic), which significantly increases memory and compute demands.
Group Relative Policy Optimization (GRPO) (Shao et al., 2024) was introduced to remove the need for a critic by computing group-level statistics over generated samples. Although effective, GRPO still assumes access to full-model fine-tuning and typically requires the storage of full trajectories for reward normalization, limiting its applicability in low-resource settings.
@@ 2.2 Memory-Efficient Fine-Tuning
Parameter-efficient fine-tuning methods such as Low-Rank Adaptation (LoRA) (Hu et al., 2021) have enabled researchers to adapt large models on modest hardware by injecting trainable rank-decomposed matrices into frozen weights. These techniques allow significant reductions in memory usage and training time without substantially sacrificing downstream performance. Our work builds on this line of research by evaluating RL algorithms in the context of LoRA-based fine-tuning, rather than full backpropagation over all model weights.
@@ 2.3 Reasoning Benchmarks
Benchmarks such as GSM8K (Cobbe et al., 2021) and SVAMP (Patel et al., 2021) have been widely adopted to measure mathematical and symbolic reasoning in language models. SVAMP in particular requires verbal reasoning over numerical quantities and has been shown to be out-of-distribution (OOD) for many LLMs (such as by presenting irrelevant information in questions) (Patel et al., 2021).
@@ 3.1 Problem Setup
Let $\pi_{\theta}(y_{t}\mid y_{<t},X)$ denote the autoregressive policy defined by the language model with parameters $\theta$. The objective is to optimize $\theta$ to maximize the expected reward over the distribution of completions:
@@ 3.2 S-GRPO: Stochastic Group Relative Policy Optimization
$\mathcal{J}_{\text{SGRPO}}(\theta)$ $=\frac{1}{|\mathcal{G}|}\sum_{i=1}^{\mathcal{G}}\frac{1}{|\mathcal{T}_{i}|}\sum_{t\in\mathcal{T}_{i}}\frac{\pi_{\theta}(y^{(i)}_{t}\mid x,y^{(i)}_{<t})}{\pi_{\theta}(y^{(i)}_{t}\mid x,y^{(i)}_{<t})\big{|}_{\text{no grad}}}\hat{A}_{i,t}$
$\quad-\,\beta\,D_{\text{KL}}\!\left[\pi_{\theta}\;\|\;\pi_{\text{ref}}\right]$ (5)
@@ 3.3 T-SPMO: Token-Specific Prefix Matching Optimization
Token-Specific Prefix Matching Optimization (T-SPMO) is a token-level RL algorithm that enables fine-grained credit assignment without trajectory-wide statistics. Similarly to GRPO, we sample $|\mathcal{G}|$ completions per prompt. From these completions, we construct a prefix tree to identify all unique $(p,v)$ token transitions, where $p$ is a prefix of tokens and $v$ is the next token.
@@
As most generations will diverge from one another near the start of the generation, we introduce a configurable replay mechanic that serves to build prefix trees in later sections of completions. This allows the model to learn strategies past the beginning of generation.
@@ 4.1 Experiment Setup
We evaluate our RL methods on two reasoning benchmarks: the SVAMP dataset and a custom multi-digit multiplication task. Each dataset is handled as a separate training run, with distinct hyperparameter configurations tuned for the nature of the task.
@@ 4.3 Ablations
Across SVAMP, S-GRPO was relatively robust to changes in $\alpha$ and $k$, with accuracies remaining between 66% and 70.3%.
@@
We also ablated the effective batch size for S-GRPO by accumulating gradients across multiple prompts before each optimization step (with total problems seen held constant):
@@
Increasing the effective batch size to 8 led to a significant degradation in performance across both tasks, likely because updates to the policy occur less frequently.
Finally, for T-SPMO, we ablated the group size $|\mathcal{G}|$ and whether completions were replayed based on success or failure:
@@
T-SPMO showed strong performance on SVAMP across all configurations. The multiplication benchmark proved to be more sensitive to hyper-parameters. Decreasing $|\mathcal{G}|$ consistently reducing performance on multiplication, likely due to insufficient exploration of alternative token transitions during optimization. Running T-SPMO without replay from a successful completion also led to significant performance degradation on multiplication, likely because T-SPMO does not optimize the model for later tokens in completions. Interestingly, for multiplication, including failed completions for replay degraded performance, indicating that forcing the model to recover from early errors may inject too much variance into training.

### v3

@@ Abstract
We propose reinforcement learning (RL) strategies tailored for reasoning in large language models (LLMs) under strict memory and compute limits, with a particular focus on compatibility with LoRA fine-tuning. Building on early policy gradient methods with baseline subtraction, we design critic-free methods that operate on a small, informative subset of output tokens to reduce memory usage and stabilize training. We introduce S-GRPO, a stochastic variant of Group Relative Policy Optimization, and T-SPMO, a token-level prefix matching approach for fine-grained credit assignment. Applied to Qwen2-1.5B, our methods raise accuracy on the SVAMP benchmark from 46% to over 70% and show strong performance on multi-digit multiplication. Surprisingly, full-token GRPO under LoRA fails to improve over the base model, suggesting that selective token-level optimization may act as an implicit regularizer in low-parameter training regimes.
@@ 1 Introduction
In this work, we design and evaluate RL algorithms that are explicitly tailored to low compute constraints: Partial parameter updates, limited GPU capacity, and the need to mitigate overfitting in low-capacity adaptation settings. First, we introduce Stochastic-GRPO (S-GRPO), a lightweight variant of GRPO that samples tokens from output trajectories to contribute to the loss. Second, we propose Token-Specific Prefix Matching Optimization (T-SPMO), which assigns credit at a token granularity. Our methods extend early baseline-based policy gradients (e.g., Kool et al., 2019; Williams, 1992) to a fine-grained token-level setting under LoRA constraints.
To evaluate these methods, we fine-tune the open-weight Qwen2-1.5B model (Yang et al., 2024) using LoRA (Hu et al., 2021) on two reasoning benchmarks: the SVAMP dataset (Patel et al., 2021) and a multi-digit multiplication task. Despite updating the policy using only 30–50% of generated tokens in S-GRPO and under 5% in T-SPMO, both methods substantially improve test-set accuracy across both benchmarks. We also establish a full-token GRPO baseline under LoRA, which surprisingly did not improve performance from the base model. These results suggest that carefully designed, memory-efficient RL algorithms can significantly improve LLM reasoning under compute constraints—even outperforming exhaustive full-token optimization strategies.
@@ 2 Related Work
2.1 RL for Language Models
Reinforcement learning has been widely used to align language model outputs with task-specific or user-defined objectives. Proximal Policy Optimization (PPO) (Schulman et al., 2017) is a standard RLHF method but requires a separate value network, increasing memory demands. Group Relative Policy Optimization (GRPO) (Shao et al., 2024) removes the critic by using group-level reward normalization, yet still relies on full-model fine-tuning and full-sequence loss computation, limiting its applicability under compute constraints.
2.2 Baselines and Grouped Advantage Estimation
Classic policy gradient methods reduce variance by subtracting a baseline from the return (Williams, 1992). Later work (Kool et al., 2019) demonstrated that using average rewards over sampled completions—rather than a learned value function—can yield strong performance. We extend this idea to token-level credit assignment for LLMs.
@@ 2.3 Memory-Efficient Fine-Tuning
Techniques like Low-Rank Adaptation (LoRA) (Hu et al., 2021) enable parameter-efficient adaptation of large models on modest hardware by training low-rank adapters in frozen networks. We focus on RL methods that are compatible with LoRA.
@@ 2.4 Reasoning Benchmarks
Datasets like GSM8K (Cobbe et al., 2021) and SVAMP (Patel et al., 2021) benchmark symbolic and arithmetic reasoning in LLMs. SVAMP poses particular challenges due to its out-of-distribution phrasing and distractor patterns, making it a strong testbed for evaluating reasoning-focused RL methods.
@@ 3.1 Problem Setup
Let $\pi_{\theta}(y_{t}\mid x,y_{<t})$ denote the autoregressive policy defined by the language model with parameters $\theta$. The objective is to optimize $\theta$ to maximize the expected reward over the distribution of completions:
@@ 3.2 S-GRPO: Stochastic Group Relative Policy Optimization
$\mathcal{J}_{\text{SGRPO}}(\theta)$ $=\frac{1}{|\mathcal{G}|}\sum_{i=1}^{\mathcal{G}}\frac{1}{|\mathcal{T}_{i}|}\sum_{t\in\mathcal{T}_{i}}(\frac{\pi_{\theta}(y^{(i)}_{t}\mid x,y^{(i)}_{<t})}{\pi_{\theta}(y^{(i)}_{t}\mid x,y^{(i)}_{<t})\big{|}_{\text{no grad}}}\hat{A}_{i,t}$
$\quad-\,\beta\,D_{\text{KL}}\!\left[\pi_{\theta}\;\|\;\pi_{\text{ref}}\right])$ (5)
@@ 3.3 T-SPMO: Token-Specific Prefix Matching Optimization
Token-Specific Prefix Matching Optimization (T-SPMO) is a token-level RL algorithm that enables fine-grained credit assignment without trajectory-wide statistics. Similarly to GRPO, we sample $|\mathcal{G}|$ completions per prompt. Inspired by grouped baselines in REINFORCE (Kool et al., 2019; Williams, 1992), we apply advantage estimation to token-level prefix transitions. We construct a token prefix trie from $\mathcal{G}$ and find all unique $(p,v)$ token transitions, where $p$ is a prefix of tokens and $v$ is the next token.
@@
As most generations will diverge from one another near the start of the generation, we introduce a configurable replay mechanic that serves to build prefix tries in later sections of completions. This allows the model to learn strategies past the beginning of generation.
@@ 4.1 Experiment Setup
We evaluate our RL methods on two reasoning benchmarks: the SVAMP dataset and a custom arithmetic task consisting of multiplication of two 3 digit numbers. Each dataset is handled as a separate training run, with distinct hyperparameter configurations tuned for the nature of the task.
@@ 4.2 Results
Table 3: Average number of tokens per example contributing to the loss during the first 100 epochs of RL training for GRPO, S-GRPO, and T-SPMO. Includes replay tokens for T-SPMO.
Method SVAMP Mult
GRPO 1204.5 2166.8
S-GRPO 557.8 674.6
T-SPMO 21.0 54.1
Table 3 reports the average number of loss-contributing tokens per example for each method, measured over the first 100 epochs of training. As the model converges, its outputs become more stable and deterministic, leading to fewer loss tokens—especially for methods like T-SPMO. We therefore focus on early training as a practical upper bound on memory usage. Both S-GRPO and T-SPMO substantially reduce the number of optimized tokens, using less than 50% and 5%, respectively, of what GRPO consumes.
@@ 4.3 Ablations
Table 4 shows that across SVAMP, S-GRPO was relatively robust to changes in $\alpha$ and $k$, with accuracies remaining between 66% and 70.3%.
@@
We also ablated the effective batch size for S-GRPO by accumulating gradients across multiple prompts before each optimization step (with total problems seen held constant).
@@
Table 5 shows that increasing the effective batch size to 8 led to a significant degradation in performance across both tasks, likely because updates to the policy occur less frequently.
Finally, for T-SPMO, we ablated the group size $|\mathcal{G}|$ and whether completions were replayed based on success or failure.
@@
We see in Table 6 that T-SPMO showed strong performance on SVAMP across all configurations. The multiplication benchmark proved to be more sensitive to hyper-parameters. Decreasing $|\mathcal{G}|$ consistently reducing performance on multiplication, likely due to insufficient exploration of alternative token transitions during optimization. Running T-SPMO without replay from a successful completion also led to significant performance degradation on multiplication, likely because T-SPMO does not optimize the model for later tokens in completions. Interestingly, for multiplication, including failed completions for replay degraded performance, indicating that forcing the model to recover from early errors may inject too much variance into training.